# Generalized Neural Sorting Networks with Error-Free Differentiable Swap Functions

### v1

@@ 4 Neural Sorting Networks with Error-Free Differentiable Swap Functions
where $\mathbf{P}_{\mathrm{gt}}$ is a ground-truth permutation matrix. Note that all the operations in $\mathcal{L}_{\mathrm{soft}}$ are entry-wise. Similar to (17), the objective (18) for $\mathbf{P}_{\mathrm{hard}}$ might be designed as the form of binary cross-entropy, which tends to be generally robust in training deep neural networks. However, we struggle to apply the binary cross-entropy for $\mathbf{P}_{\mathrm{hard}}$ into our problem formulation, due to discretized loss values. In particular, the form of cross-entropy for $\mathbf{P}_{\mathrm{hard}}$ can be used to train the sorting network, but degrades its performance in our preliminary experiments. Thus, we design the objective for $\mathbf{P}_{\mathrm{hard}}$ as $\mathcal{L}_{\mathrm{hard}}=\|\mathbf{X}_{\mathrm{o},\mathrm{hard}}-\mathbf{X}_{\mathrm{o},\mathrm{gt}}\|_{F}^{2}$ with the Frobenius norm, which helps to train the network more robustly.
In addition, using a proposition on splitting $\mathbf{P}_{\mathrm{hard}}$, which is discussed in Section G, the objective (18) for $\mathbf{P}_{\mathrm{hard}}$ can be modified by splitting $\mathbf{P}_{\mathrm{hard}}$, $\mathbf{P}_{\mathrm{gt}}$, and $\mathbf{X}$, which is able to reduce the number of possible permutations; see the associated section for the details. Eventually, our network $g$ is trained by the combined loss $\mathcal{L}=\mathcal{L}_{\mathrm{soft}}+\lambda\mathcal{L}_{\mathrm{hard}}$ where $\lambda$ is a balancing hyperparameter; analysis on $\lambda$ can be found in the appendices. As mentioned above, a landscape of $\mathcal{L}_{\mathrm{hard}}$ is not smooth due to the property of straight-through estimator, even though we use $\mathcal{L}_{\mathrm{hard}}$. Thus, we combine both objectives to the form of a single loss, which is widely adopted in a deep learning community.
5 Experimental Results
@@
where $\mathrm{argsort}$ returns the indices to sort a given vector and $\boldsymbol{1}(\cdot)$ is an indicator function. Note that
@@
We attempt to match the capacities of the Transformer-based models to the conventional CNNs. As described in Tables 12, and 3, the capacities of the small Transformer-based models are smaller than or similar to the capacities of the CNNs in terms of FLOPs and the number of parameters.
@@ 5.1 Sorting Multi-Digit Images
Ours Error-Free DSFs CNN 94.8 (96.4) 86.9 (94.1) 74.2 (90.9) 62.6 (88.6) 34.7 (83.3) 2.1 (69.2) 130M 855K
@@
We conduct the experiments 5 times by varying random seeds to report the average of $\mathrm{acc}_{\mathrm{em}}$ and $\mathrm{acc}_{\mathrm{ew}}$, and use an optimal monotonic sigmoid function as DSFs. The performance of each model is measured by a test dataset. We use the AdamW optimizer (Loshchilov & Hutter, 2018), and train each model for 200,000 steps on the four-digit MNIST dataset and 300,000 steps on the SVHN dataset. Unless otherwise noted, we follow the same settings of the work (Petersen et al., 2022) for fair comparisons. The missing details are described in Section I.
@@ Results.
Tables 1 and 2 show the results of the previous work such as NeuralSort (Grover et al., 2019), Sinkhorn Sort (Cuturi et al., 2019), Fast Sort & Rank (Blondel et al., 2020), and Diffsort (Petersen et al., 2021; 2022), and our methods on the MNIST and SVHN datasets, respectively. When we use the conventional CNN as a permutation-equivariant network, our method shows better than or comparable to the previous methods. As we exploit more powerful models, i.e., small and large Transformer-based permutation-equivariant models, our approaches show better results compared to other existing methods including our method with the conventional CNN. [^1] Thanks to many open-source projects, we can easily run the baseline methods. However, it is difficult to reproduce some results due to unknown random seeds. For this reason, we bring the results from the work (Petersen et al., 2022), and use fixed random seeds, i.e., 42, 84, 126, 168, 210, for our methods.
@@ 5.2 Sorting Image Fragments
Table 3: Results on sorting image fragments of MNIST and CIFAR-10. 2 $\times$ 2 and 3 $\times$ 3 indicate the number of fragments, and 14 $\times$ 14, 9 $\times$ 9, 16 $\times$ 16, and 10 $\times$ 10 (in parentheses) indicate the size of image fragments. FLOPs is computed on the basis of the MNIST 2 $\times$ 2 (14 $\times$ 14) case and the CIFAR-10 2 $\times$ 2 (16 $\times$ 16) case. All the values are averaged over 5 runs with different seeds.
@@
When inputs are fragments, we use two datasets: the MNIST dataset (LeCun et al., 1998) and the CIFAR-10 dataset (Krizhevsky & Hinton, 2009). Similar to the work (Mena et al., 2018), we create multiple fragments (or patches) from a single-digit image of the MNIST dataset to utilize themselves as inputs – for example, 4 fragments of size 14 $\times$ 14 or 9 fragments of size 9 $\times$ 9 are created from a single image. Similarly, the CIFAR-10 dataset, which contains various objects, e.g., birds and cats, is split to multiple patches, and then is used to the experiments on sorting image fragments. See Table 3 for the details of the fragments.
@@
Similar to the experiments on sorting multi-digit images, an optimal monotonic sigmoid function is used as DSFs. Since the size of inputs is much smaller than the experiments on sorting multi-digit images, shown in Section 5.1, we modify the architectures of the CNNs and the Transformer-based models. We reduce the kernel size of convolution layers from five to three and make strides two. Due to the small input size, we omit the results by a large Transformer-based model for these experiments. Additionally, max-pooling operations are removed. Similar to the experiments in Section 5.1, we use the AdamW optimizer (Loshchilov & Hutter, 2018). Moreover, each model is trained for 50,000 steps when the number of fragments is $2\times 2$, i.e., when a sequence length is 4, and 100,000 steps for $3\times 3$ fragments, i.e., when a sequence length is 9. Additional information including the details of neural architectures can be found in the appendices.
@@ Results.
According to the experimental results, we achieve satisfactory performance by applying error-free DSFs, combined loss, and Transformer-based models with multi-head attention. We provide detailed discussion on how they contribute to the performance gain in Section 7, and empirical studies on steepness, learning rate, and a balancing hyperparameter in Section 7 and the appendices.
@@ 6 Related Work
To allow us to differentiate a sorting algorithm, Grover et al. (2019) have proposed the continuous relaxation of $\mathrm{argsort}$ operator, which is named NeuralSort. In this work the output of NeuralSort only satisfies the row-stochastic matrix property, although Grover et al. (2019) attempt to employ a gradient-based optimization strategy in learning a neural sorting algorithm. Cuturi et al. (2019) propose the smoothed ranking and sorting operators using optimal transport, which is the natural relaxation for assignments. To reduce the cost of the optimal transport, the Sinkhorn algorithm (Cuturi, 2013) is used. Then, the reference (Blondel et al., 2020) has proposed the first differentiable sorting and ranking operators with $\mathcal{O}(n\log n)$ time and $\mathcal{O}(n)$ space complexities, which is named Fast Rank & Sort, by constructing differentiable operators as projections on permutahedron. The work (Petersen et al., 2021) has suggested a differentiable sorting network with relaxed conditional swap functions. Recently, the same authors analyze the characteristics of the relaxation of monotonic conditional swap functions, and propose several monotonic swap functions, e.g., Cauchy and optimal monotonic functions (Petersen et al., 2022).
@@
A seminal architecture, long short-term memory (Hochreiter & Schmidhuber, 1997) can be used in modeling a sequence without any difficulty, and a sequence-to-sequence model (Sutskever et al., 2014) can be employed to cope with a sequence. However, as discussed in the work (Vinyals et al., 2016), an unordered sequence can have good orderings, by analyzing the effects of permutation thoroughly. Zaheer et al. (2017) propose a permutation-invariant or permutation-equivariant network, named Deep Sets, and prove the permutation invariance and permutation equivariance of the proposed models. By utilizing a Transformer network (Vaswani et al., 2017), Lee et al. (2019) propose a permutation-equivariant network.
@@ 7 Discussion
Table 4: Study on comparisons of Diffsort with an optimal monotonic sigmoid function and our methods in the experiments of sorting the four-digit MNIST dataset
@@
We carry out a numerical analysis on the effects of our methods, compared to a baseline method, i.e., Diffsort with an optimal monotonic sigmoid function. As presented in Table 4, we demonstrate that our methods better sorting performance compared to the baseline, which implies that our suggestions are effective in the sorting tasks. In these experiments, we follow the settings of the experiments described in Section 5.1. Moreover, we demonstrate numerical analyses on a balancing hyperparameter, steepness, and learning rate in the appendices.
@@
According to the results in Sections 5 and 7, we can argue that error-free DSFs, our proposed loss, and Transformer-based models contribute to better performance appropriately compared to the baseline methods. As shown in Tables 1 and 4, the performance gains by Transformer-based models are more substantial than the gains by error-free DSFs and our loss, since multi-head attention is effective for capturing long-term dependency (or dependency between multiple instances in our case) and reducing inductive biases. However, as will be discussed in the following, the hard permutation matrices can be used in the case that does not allow us to mix instances in $\mathbf{X}$, e.g., sorting image fragments in Section 5.2.
@@
While a soft permutation matrix $\mathbf{P}_{\mathrm{soft}}$ is limited to be utilized itself directly, a hard permutation matrix $\mathbf{P}_{\mathrm{hard}}$ is capable of being applied in a case that requires $\mathbf{P}_{\mathrm{hard}}$; each row of $\mathbf{P}_{\mathrm{soft}}^{\top}\mathbf{X}$ is a linear combination of some column of $\mathbf{P}_{\mathrm{soft}}$ and $\mathbf{X}$, but one of $\mathbf{P}_{\mathrm{hard}}^{\top}\mathbf{X}$ corresponds to an exact row in $\mathbf{X}$. The experiments in Section 5.2 are one of such cases, and it exhibits our strength, not only the performance in $\mathrm{acc}_{\mathrm{em}}$ and $\mathrm{acc}_{\mathrm{ew}}$.
@@
We follow the model architecture, which is used in the previous work (Grover et al., 2019; Cuturi et al., 2019; Petersen et al., 2021; 2022), for the CNNs. However, as shown in Tables 12, and 3, the model is not enough to show the best performance. In particular, whereas the model capacity, i.e., FLOPs and the number of parameters, of the small Transformer models is almost matched to or less than the capacity of the CNNs, the results by the small Transformer models outperform the results by the CNNs. We presume that these performance gains are derived from a multi-head attention’s ability to capture long-term dependency and reduce an inductive bias, as widely stated in many recent studies in diverse fields such as natural language processing (Vaswani et al., 2017; Devlin et al., 2018; Brown et al., 2020), computer vision (Dosovitskiy et al., 2021; Liu et al., 2021), and 3D vision (Nash et al., 2020; Zhao et al., 2021). Especially, unlike the instance-wise CNNs, our permutation-equivariant Transformer architecture utilizes self-attention given instances, so that our model can productively compare instances in a sequence and effectively learn the relative relationship between them.
@@ 8 Conclusion
In this paper, we defined a softening error, induced by a monotonic DSF, and demonstrated several evidences of the error accumulation problem. To resolve the error accumulation problem, an error-free DSF is proposed, inspired by a straight-through estimator. Moreover, we provided the simple theoretical and empirical analyses that our error-free DSF successfully achieves a zero error and also holds non-decreasing and differentiability conditions. By combining all components, we suggested a generalized neural sorting network with the error-free DSF using multi-head attention. Finally we showed that our methods are better than or comparable to other algorithms in diverse benchmarks.
Appendix A Optimal Monotonic Sigmoid Function
Figure 4: An optimal monotonic sigmoid function, shown in (7)
@@ Appendix A Optimal Monotonic Sigmoid Function
Appendix B Visualization of Sorting Network
Figure 5: A sorting network with 5 wire sets
Figure 5 shows an example of a sorting network and permutation matrices.
Appendix C Comparisons of Differentiable Swap Functions
Figure 6: Comparisons of diverse DSFs in terms of how many times a swap function is conducted. Our error-free DSF does not change the original $x$ and $y$, compared to other DSFs. We set (left) $x=4,y=0$ or (right) $x=8,y=0$ initially. We conduct a swap functions $k$ times where $k=\#\textrm{Swaps}$.
As shown in Figure 6, some sigmoid functions, i.e., logistic, logistic with ART, reciprocal, Cauchy, and optimal monotonic functions suffer from the error accumulation problem; see the work (Petersen et al., 2022) for the details of such sigmoid functions. For the case of the Cauchy function, two values are close enough at $9$th step in the left panel of Figure 6 and $15$th step in the right panel of Figure 6; we present the corresponding step where a difference between two values becomes smaller than 0.001.

### v2

@@ 4 Neural Sorting Networks with Error-Free Differentiable Swap Functions
where $\mathbf{P}_{\mathrm{gt}}$ is a ground-truth permutation matrix. Note that all the operations in $\mathcal{L}_{\mathrm{soft}}$ are entry-wise. Similar to (17), the objective (18) for $\mathbf{P}_{\mathrm{hard}}$ should be designed as the form of binary cross-entropy, which tends to be generally robust for training deep neural networks. However, we struggle to apply the binary cross-entropy for $\mathbf{P}_{\mathrm{hard}}$ into our problem formulation, due to discretized loss values. In particular, the form of cross-entropy for $\mathbf{P}_{\mathrm{hard}}$ can be used to train the sorting network, but degrades its performance in our preliminary experiments. Thus, we choose the objective for $\mathbf{P}_{\mathrm{hard}}$ as $\|\mathbf{X}_{\mathrm{o},\mathrm{hard}}-\mathbf{X}_{\mathrm{o},\mathrm{gt}}\|_{F}^{2}$ with the Frobenius norm, which helps to train the network more robustly.
In addition, using a proposition on splitting $\mathbf{P}_{\mathrm{hard}}$, which is discussed in Section H, the objective (18) for $\mathbf{P}_{\mathrm{hard}}$ can be modified by splitting $\mathbf{P}_{\mathrm{hard}}$, $\mathbf{P}_{\mathrm{gt}}$, and $\mathbf{X}$, which is able to reduce the number of possible permutations; see the associated section for details. Eventually, our network $g$ is trained by the combined loss $\mathcal{L}=\mathcal{L}_{\mathrm{soft}}+\lambda\mathcal{L}_{\mathrm{hard}}$, where $\lambda$ is a balancing hyperparameter; an analysis on $\lambda$ can be found in the appendices. As mentioned above, a landscape of $\mathcal{L}_{\mathrm{hard}}$ is not smooth due to the property of a straight-through estimator, even though we use $\mathcal{L}_{\mathrm{hard}}$. Thus, we combine both objectives to the form of a single loss, which is widely adopted in the deep learning community.
5 Experiments
@@
where $\mathrm{argsort}$ returns indices to sort a given vector and $\boldsymbol{1}(\cdot)$ is an indicator function. Note that
@@
We attempt to match the capacities of the Transformer-based models to the conventional CNNs. As described in Tables 12, and 3, the capacities of the Transformer-Small models are smaller than or similar to the capacities of the CNNs in terms of FLOPs and the number of parameters.
@@ 5.1 Sorting Multi-Digit Images
Ours Error-Free DSFs CNN 95.2 (96.7) 87.2 (94.2) 76.6 (91.6) 64.8 (89.2) 34.7 (83.3) 2.1 (69.2) 130M 855K
@@
We conduct the experiments 5 times by varying random seeds to report the average of $\mathrm{acc}_{\mathrm{em}}$ and $\mathrm{acc}_{\mathrm{ew}}$, and use the optimal monotonic sigmoid function as DSFs. The performance of each model is measured by a test dataset. We use the AdamW optimizer (Loshchilov & Hutter, 2018), and train each model for 200,000 steps on the four-digit MNIST dataset and 300,000 steps on the SVHN dataset. Unless otherwise noted, we follow the same settings of the work (Petersen et al., 2022) for fair comparisons. Missing details are described in Section J.
@@ Results.
Tables 1 and 2 show the results of the previous work such as NeuralSort (Grover et al., 2019), Sinkhorn Sort (Cuturi et al., 2019), Fast Sort & Rank (Blondel et al., 2020), and Diffsort (Petersen et al., 2021; 2022), and our methods on the MNIST and SVHN datasets, respectively. When we use the conventional CNN as a permutation-equivariant network, our method shows better than or comparable to the previous methods. As we exploit more powerful models, i.e., the Transformer-Small and Transformer-Large permutation-equivariant models, our approaches show better results compared to other existing methods including our method with the conventional CNN. [^1] Thanks to many open-source projects, we can easily run the baseline methods. However, it is difficult to reproduce some results due to unknown random seeds. For this reason, we bring the results from the work (Petersen et al., 2022), and use fixed random seeds, i.e., 42, 84, 126, 168, 210, for our methods.
@@ 5.2 Sorting Image Fragments
Table 3: Results on sorting image fragments of MNIST and CIFAR-10. 2 $\times$ 2 and 3 $\times$ 3 indicate the numbers of fragments, and 14 $\times$ 14, 9 $\times$ 9, 16 $\times$ 16, and 10 $\times$ 10 (in parentheses) indicate the sizes of image fragments. FLOPs is computed on the basis of the MNIST 2 $\times$ 2 (14 $\times$ 14) case and the CIFAR-10 2 $\times$ 2 (16 $\times$ 16) case. All the values are averaged over 5 runs with different seeds.
@@
For experiments on sorting image fragments, we use two datasets: the MNIST dataset (LeCun et al., 1998) and the CIFAR-10 dataset (Krizhevsky & Hinton, 2009). Similar to the work (Mena et al., 2018), we create multiple fragments or patches from a single-digit image of the MNIST dataset to utilize themselves as inputs – for example, 4 fragments of size 14 $\times$ 14 or 9 fragments of size 9 $\times$ 9 are created from a single image. Similarly, the image included in the CIFAR-10 dataset, which contains one of various objects, e.g., birds and cats, is split to multiple patches, and then is used to the experiments on sorting image fragments. See Table 3 for the details of the image fragments and their sizes.
@@
Similar to the experiments on sorting multi-digit images, an optimal monotonic sigmoid function is used as DSFs. Since the size of inputs is much smaller than the experiments on sorting multi-digit images, shown in Section 5.1, we modify the architectures of the CNNs and the Transformer-based models. We reduce the kernel size of convolution layers from 5 to 3 and make strides 2. Due to the small input sizes, we omit the results by the Transformer-Large model for these experiments. Additionally, max-pooling operations are removed. Similar to the experiments in Section 5.1, we use the AdamW optimizer (Loshchilov & Hutter, 2018). Moreover, each model is trained for 50,000 steps when the number of fragments is $2\times 2$, i.e., when a sequence length is 4, and 100,000 steps for $3\times 3$ fragments, i.e., when a sequence length is 9. Additional information including the details of neural architectures can be found in Sections I and J.
@@ Results.
According to the experimental results, we achieve satisfactory performance by applying the error-free DSFs, combined loss, and Transformer-based models with multi-head attention. We provide detailed discussion on how they contribute to the performance gains in Section 7, and empirical studies on steepness, learning rate, and a balancing hyperparameter in Section 7 and the appendices.
@@ 6 Related Work
To allow us to differentiate a sorting algorithm, Grover et al. (2019) have proposed the continuous relaxation of $\mathrm{argsort}$ operator, which is named NeuralSort. In this work, the output of NeuralSort only satisfies the row-stochastic matrix property, although Grover et al. (2019) attempt to employ a gradient-based optimization strategy in learning a neural sorting algorithm. Cuturi et al. (2019) propose a smoothed ranking and sorting operator using optimal transport, which is the natural relaxation for assignments. To reduce the cost of the optimal transport, the Sinkhorn algorithm (Cuturi, 2013) is used. Then, Blondel et al. (2020) have proposed a differentiable sorting and ranking operator with $\mathcal{O}(n\log n)$ time and $\mathcal{O}(n)$ space complexities, which is named Fast Rank & Sort, by constructing differentiable operators as projections on permutahedron. Petersen et al. (2021) have suggested a differentiable sorting network with relaxed conditional swap functions. Recently, the same authors analyze the characteristics of the relaxation of monotonic conditional swap functions, and propose several monotonic swap functions, e.g., the Cauchy and optimal monotonic functions (Petersen et al., 2022).
@@
A seminal architecture, long short-term memory (Hochreiter & Schmidhuber, 1997) can be used in modeling a sequence without any difficulty, and a sequence-to-sequence model (Sutskever et al., 2014) can be employed to cope with a sequence. However, as discussed in the work by Vinyals et al. (2016), an unordered sequence can have good orderings, by analyzing the effects of permutation thoroughly. Zaheer et al. (2017) propose a permutation-invariant or permutation-equivariant network, named Deep Sets, and prove the permutation invariance and permutation equivariance of the proposed models. By utilizing the Transformer network (Vaswani et al., 2017), Lee et al. (2019) have proposed a permutation-equivariant network.
@@ 7 Discussion
Table 4: Study on comparisons of Diffsort with the optimal monotonic sigmoid function and our methods in the experiments on sorting the four-digit MNIST dataset.
@@
We carry out numerical analyses on the effects of our methods, compared to a baseline method, i.e., Diffsort with the optimal monotonic sigmoid function. As reported in Table 4, we demonstrate that our methods better sorting performance compared to the baseline, which implies that our suggestions are effective in the sorting tasks. In these experiments, we follow the settings of the experiments described in Section 5.1. Moreover, we present numerical analyses on a balancing hyperparameter, steepness, and a learning rate in Sections K and L.
@@
According to the results in Sections 5 and 7, we can argue that the error-free DSFs, our proposed loss, and the Transformer-based models contribute to better performance considerably compared to the baseline methods. As shown in Tables 1 and 4, the performance gains by the Transformer-based models are more substantial than the gains by the error-free DSFs and our loss, since multi-head attention is effective for capturing long-term dependency (or dependency between multiple instances in our case) and reducing inductive biases. However, as will be discussed in the following, the hard permutation matrices can be used in the case that does not allow us to mix instances in $\mathbf{X}$, e.g., sorting image fragments in Section 5.2.
@@
While the use of a soft permutation matrix $\mathbf{P}_{\mathrm{soft}}$ makes given instances mixed, a hard permutation matrix $\mathbf{P}_{\mathrm{hard}}$ is instrumental in applying $\mathbf{P}_{\mathrm{hard}}$ in a problem that requires swapping given instances exactly. More precisely, each row of $\mathbf{P}_{\mathrm{soft}}^{\top}\mathbf{X}$ is a linear combination of some column of $\mathbf{P}_{\mathrm{soft}}$ and $\mathbf{X}$, but one of $\mathbf{P}_{\mathrm{hard}}^{\top}\mathbf{X}$ corresponds to an exact row in $\mathbf{X}$. This property can be used to preserve input instances from sorting operations. The experiments in Section 5.2 can be considered as one of such cases, and it exhibits the strength of our method, not only the performance in $\mathrm{acc}_{\mathrm{em}}$ and $\mathrm{acc}_{\mathrm{ew}}$.
@@
We follow the model architecture used in the previous work (Grover et al., 2019; Cuturi et al., 2019; Petersen et al., 2021; 2022) for the CNNs. However, as shown in Tables 12, and 3, the model is not enough to show the best performance. In particular, whereas the model capacity, i.e., FLOPs and the number of parameters, of the Transformer-Small models is almost matched to or less than the capacity of the CNNs, the results by the Transformer-Small models outperform the results by the CNNs. We presume that these performance gains are derived from a multi-head attention’s ability to capture long-term dependency and reduce inductive biases, as widely stated in many recent studies in diverse fields such as natural language processing (Vaswani et al., 2017; Devlin et al., 2018; Brown et al., 2020), computer vision (Dosovitskiy et al., 2021; Liu et al., 2021), and 3D vision (Nash et al., 2020; Zhao et al., 2021). Especially, unlike the instance-wise CNNs, our permutation-equivariant Transformer architecture utilizes self-attention for given instances, so that our model can productively compare instances in a sequence and effectively learn the relative relationship between them.
@@ 8 Conclusion
In this paper, we defined a softening error, induced by a monotonic DSF, and demonstrated several evidences of the error accumulation problem. To resolve the error accumulation problem, an error-free DSF is proposed, inspired by a straight-through estimator. Moreover, we provided the simple theoretical and empirical analyses that our error-free DSF successfully achieves a zero error and also holds a non-decreasing condition and differentiability. By combining all components, we suggested a generalized neural sorting network with the error-free DSF and multi-head attention. Finally, we showed that our methods are better than or comparable to other algorithms in diverse benchmarks.
Ethics Statement
As discussed in Section 7, the hard permutation matrices produced by our methods allow us to swap instances exactly, not the linear combination of instances. This characteristic is required when we are given the final outcomes of sorting as supervision. This scenario is tested by the experiments presented in Section 5.2. In these experiments, we are supposed that original images are provided as supervision. Building on the advantages of neural network-based sorting networks, we expand their practical significance to the cases that need hard permutation matrices. On the other hand, the nature of neural sorting networks may yield a potential negative societal impact. If this line of research including our approaches is employed to sort controversial high-dimensional data such as beauty and intelligence, it can be considered as the unethical use cases of artificial intelligence.
Acknowledgments
This work was supported by the IITP grants (2022-0-00290: Visual Intelligence for Space-Time Understanding and Generation based on Multi-layered Visual Common Sense, 2022-0-00264: Comprehensive Video Understanding and Generation with Knowledge-based Deep Logic Neural Network) funded by Ministry of Science and ICT, Republic of Korea.
Appendix A Optimal Monotonic Sigmoid Functions
Figure 4: An optimal monotonic sigmoid function, which is presented in (7).
@@ Appendix A Optimal Monotonic Sigmoid Functions
Appendix B Comparisons of Differentiable Swap Functions
Figure 5: Comparisons of diverse DSFs in terms of the numbers of swap functions applied. Our error-free DSF does not change the original $x$ and $y$, unlike other DSFs. We initially set $x=4,y=0$ for the left panel or $x=8,y=0$ for the right panel, where $k=\#\textrm{Swaps}$.
As depicted in Figure 5, some sigmoid functions such as the logistic, logistic with ART, reciprocal, Cauchy, and optimal monotonic functions suffer from the error accumulation problem; see the work (Petersen et al., 2022) for the details of such sigmoid functions. For the case of the Cauchy function, two values are close enough at the $9$th step in the left panel of Figure 5 and the $15$th step in the right panel of Figure 5; we calculate the corresponding steps where a difference between two values becomes smaller than 0.001.
Appendix C Comparisons of Different Sorting Networks
Figure 6: $\mathrm{acc}_{\mathrm{em}}$ (solid) and $\mathrm{acc}_{\mathrm{ew}}$ (dashed), versus sequence lengths. Without a mapping $g$ from $\mathbf{x}$ to $s$, we uniformly sample $n$ elements from $[-10,10]$, which are considered as ordinal scores $\mathbf{s}$. Then, we sort them using the respective sorting networks and measure accuracy over the ground-truth permutation matrices computed by $n$ elements. We repeat each experiment 10,000 times.
Figure 6 shows the comparisons of different sorting networks by varying sequence lengths. $\mathrm{acc}_{\mathrm{em}}$ and $\mathrm{acc}_{\mathrm{ew}}$ are measured to assess the sorting networks.